OpenStudy (anonymous):

Write the first trigonometric function in terms of the second for θ in the given quadrant. sec θ, sin θ; θ in quadrant I

OpenStudy (anonymous):

1/sqrt(1-sin^2(theta) but how do i get this

OpenStudy (campbell_st):

so \[\sec(\theta) = \frac{1}{\cos(\theta)}\]  and \[\sin(\theta) = \cos(90 - \theta)\] by the complementary angle rule so \[\cos(\theta) = \sin(90 - \theta)\] so use this to express \[\sec(\theta)\] in terms of sin